# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 5. Necessary condition for a hyperbolic group to have a tree of graphs as its Gromov boundary
Let $G$ by a hyperbolic group and assume that $G$ is not virtually free. Then $G$ has a $1$-ended quasi-convex subgroup.
@@
By the Dunwoody’s Accessibility Theorem [12], there is a splitting of $G$ as a finite graph of groups whose edge groups are finite and whose vertex groups have at most $1$ end. Since graph of finite groups always yields a virtually free group (see Proposition 11 in [20]), at least one of the vertex groups above must be 1-ended. Finally, this vertex group is quasi-convex e.g. by the discussion in the paragraph right before Corollary 3.7. ∎
Next rather well known result follows e.g. by [14, Theorem 2.28].
@@
[…]
@@
For the sake of reaching a contradiction, assume that $\partial G$ is a tree of graphs and $G$ has a rigid cluster factor $H$ that is not virtually free. Then, by Lemma 5.9, the rigid cluster factor $H$ has a $1$-ended quasi-convex subgroup $K$. Then, since $H$ is quasi-convex in $G$, the boundary $\partial K$ is a closed subspace of $\partial G$. Since $K$ is $1$-ended, the boundary $\partial K$ is connected. By Lemma 5.10, the boundary $\partial K$ has at least two points. Moreover, the boundary $\partial G$ is connected, locally connected and has no cutpoints [14, Proposition 7.1, Theorem 7.2]. Thus we can apply Proposition 5.7 to the subspace $\partial K\subset\partial G$ to conclude that $\partial K$ has uncountable intersection with some subnecklace $N^{\prime}\subset\partial G$. Recalling Remark 5.2, we see that $\partial K$ has uncountable intersection with a necklace class $N$ of $\partial G$. Then, since $\partial K\subset\partial H$, we see that $\partial H$ has uncountable intersection with $N$. This contradicts Lemma 5.8. ∎
@@ 5.2. Twin graphs and trees of thick $\theta$-graphs
[…]
@@
Lemma 5.15 (Bowditch [1, Lemma 3.8]).
@@
We now prove (2). Suppose $\deg(y_{i})\neq\bigl{|}\pi_{0}\bigl{(}X_{i}\setminus\{x_{i},y_{i}\}\bigr{)}\bigr{|}$. Without loss of generality, we can assume $i=1$. Notice that $\deg(y_{1})$ is equal to the cardinality of the blow-up divisor $P_{y_{1}}$ and that $\bigl{|}\pi_{0}\bigl{(}X_{1}\setminus\{x_{1},y_{1}\}\bigr{)}\bigr{|}$ is equal to the number of components of the blow-up $X_{1}^{\#}(\{x_{1},y_{1}\})$. Then, by Lemma 4.6, we have distinct $p,p^{\prime}\in P_{y_{1}}$ that are contained in the same component $C$ of $X_{1}^{\#}(\{x_{1},y_{1}\})$. But $X_{1}^{\#}(x_{1})$ is a subspace of $X$ with $y_{1}\in X_{1}^{\#}(x_{1})$ and $y_{2}\notin X_{1}^{\#}(x_{1})$ so $X_{1}^{\#}(\{x_{1},y_{1}\})$ is a subspace of $X^{\#}(\{y_{1},y_{2}\})$. Then $C$ is a connected subspace of $X^{\#}(\{y_{1},y_{2}\})$ containing $p,p^{\prime}\in P_{y_{1}}$, which by Remark 5.13, implies that $y_{1}$ and $y_{2}$ are not twins.
Suppose now that $\ell(\sigma_{1})\neq\sigma_{2}$. Without loss of generality (by swapping, if necessary, the indices $i\in\{1,2\}$) there are $p,p^{\prime}\in P_{x_{1}}$ contained in the same component $C_{1}$ of $X_{1}^{\#}(\{x_{1},y_{1}\})$ such that $\ell(p)$ and $\ell(p^{\prime})$ are contained in distinct components $C_{2}$ and $C_{2}^{\prime}$ of $X_{2}^{\#}(\{x_{2},y_{2}\})$. By Lemma 4.6, there are $p_{2},p_{2}^{\prime}\in P_{y_{2}}$ with $p_{2}\in C_{2}$ and $p_{2}^{\prime}\in C_{2}^{\prime}$. We can view $X^{\#}(\{y_{1},y_{2}\})$ as the connected sum of $X_{1}^{\#}(\{x_{1},y_{1}\})$ and $X_{2}^{\#}(\{x_{2},y_{2}\})$ along $\ell\colon P_{x_{1}}\to P_{x_{2}}$. Then $C_{1}\cup C_{2}\cup C_{2}^{\prime}$ is a connected subspace of $X^{\#}(\{y_{1},y_{2}\})$ containing both $p_{2}$ and $p_{2}^{\prime}$ which, by Remark 5.13, implies that $y_{1}$ and $y_{2}$ are not twins. Thus we have proved the forward implication of (2).
@@
Lemma 5.22 (Bowditch [1, Lemma 3.18]).
@@
For an essential twin pair $\{x,\bar{x}\}$, let $n_{\{x,\bar{x}\}}$ be the maximal number of vertices contained in a component of $X\setminus\{x,\bar{x}\}$. The complexity of $\{x,\bar{x}\}$ is the nonnegative integer $c_{\{x,\bar{x}\}}=n-n_{\{x,\bar{x}\}}-2$, where $n$ is the number of essential vertices of $X$. Note that $c_{\{x,\bar{x}\}}$ is the number of vertices in all but a “largest” component of $X\setminus\{x,\bar{x}\}$. So it suffices to find an essential twin pair $\{x,\bar{x}\}$ for which $c_{\{x,\bar{x}\}}=0$. This is the base case of an induction starting with an arbitrary essential twin pair $\{x,\bar{x}\}$. It suffices to show that if $c_{\{x,\bar{x}\}}>0$ then we can find an essential twin pair $\{y,\bar{y}\}$ for which $c_{\{y,\bar{y}\}}<c_{\{x,\bar{x}\}}$. Let $C^{\prime}$ be a component of $X\setminus\{x,\bar{x}\}$ containing $n_{\{x,\bar{x}\}}$ essential vertices. Since $n_{\{x,\bar{x}\}}=n-c_{\{x,\bar{x}\}}-2<n-2$, some other component $C$ of $X\setminus\{x,\bar{x}\}$ contains an essential vertex $y$. We will show that $c_{\{y,\bar{y}\}}<c_{\{x,\bar{x}\}}$.
@@
Suppose $C$ contains at least one essential vertex. By Lemma 4.6 there is a topological edge $e_{x}$ with endpoint $x$ and whose interior is contained in $C$. Similarly, there is a topological edge $e_{\bar{x}}$ with endpoint $\bar{x}$ and whose interior is contained in $C$. Let $p_{x}$ be a point in the interior of $e_{x}$ and let $p_{\bar{x}}$ be a point in the interior of $e_{\bar{x}}$. Then $X^{\#}(p_{x},p_{\bar{x}})$ has two components, one of which contains $x$ and $\bar{x}$ and the other of which is naturally a subspace of $C$. Gluing together the pairs of points of the blow-up divisors $P_{x}\cup P_{\bar{x}}\subset X^{\#}(p_{x},p_{\bar{x}})$ that are contained in the same component results in a disjoint union of two graphs: a thick $\theta$-graph $X_{1}$ with essential vertices $x$ and $\bar{x}$ and a graph $X_{2}$ with two less essential vertices than $X$. Moreover, the original graph $X$ is a connected sum of $X_{1}$ and $X_{2}$. Thus, by induction, it suffices to show that $X_{2}$ is a twin graph. This follows immediately from Proposition 5.19. ∎
@@
In the inductive step, we can write $X$ as a connected sum $X_{1}\#X_{2}$ where $X_{1}$ is a twin graph and $X_{2}$ is a thick $\theta$-graph and the connected sum is performed at points $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Since performing a connected sum with a $\theta$-graph at an essential vertex does not change the topology, we can assume that $\deg(x_{1})=\deg(x_{2})=2$. Up to homeomorphism, we can assume that $x$ and $y$ are not contained in the blow-up divisor $P_{x_{1}}=P_{x_{2}}\subset X_{1}\#X_{2}$. In fact, we can assume that $x$ and $y$ are not contained in $e\setminus\{v,\bar{v},x_{2}\}\subset X_{1}\#X_{2}$ where $e$ is the topological edge of $X_{2}$ containing $x_{2}$ and $\{v,\bar{v}\}$ is the unique essential twin pair of $X_{2}$. We consider first the case where one of the points, say $x$, is contained in $X_{2}\setminus e\subset X_{1}\#X_{2}$. Then $y$ cannot be contained in the topological edge of $X_{2}$ that contains $x$ since otherwise $X\setminus\{x,y\}$ would be disconnected. Hence $x$ and $y$ are contained in distinct components of the complement of $\{v,\bar{v}\}$, which is a twin pair of $X$ by Proposition 5.19(1). It remains then to consider the case where $x,y\in X_{1}\setminus\{x_{1}\}\subset X_{1}\#X_{2}$. By Proposition 5.19(1), the degree $2$ points $x$ and $y$ are not twins in $X_{1}$ and so $X_{1}\setminus\{x,y\}$ is connected. Then, by induction, there exists an essential twin pair $\{w,\bar{w}\}$ of $X_{1}$ separating $x$ from $y$. By Proposition 5.19(1), the pair $\{w,\bar{w}\}$ remain twins in $X_{1}\#X_{2}$. Since the connected sum operation with the $\theta$-graph $X_{2}$ preserves components of the complement of $\{w,\bar{w}\}$ we see that $\{w,\bar{w}\}$ continues to separate $x$ and $y$ in $X=X_{1}\#X_{2}$. ∎
[…]
@@
We first prove the contrapositive of the forward implication. Assume that for any finite b-subtree $F_{0}\subset T$ there exists a finite b-subtree $F\supset F_{0}$ such that the projection $x_{F}$ of $x$ to the reduced partial union $K_{F}^{\ast}$ has degree less than $3$. Let $U$ be an open neighborhood of $x$ in $Z$. Then there exists an open neighborhood $V_{F}$ of the projection $x_{F}$ of $x$ in a reduced partial union $K_{F}^{\ast}$ such that the preimage of $V_{F}$ in $Z$ is contained in $U$. By our assumption, we can choose $K_{F}^{\ast}$ so that $x_{F}$ has degree $2$ and so we can choose $V_{F}$ so that it is homeomorphic to an open interval. By Lemma 4.13, the preimage $V$ of $V_{F}$ in $Z$ is connected. It suffices to show that $V\setminus\{x\}$ has at most two components in order to establish that $x$ has degree at most $2$. Now, let $K_{\hat{F}}^{\ast}$ be any reduced partial union with $\hat{F}\supset F$, let $V_{\hat{F}}$ be the preimage of $V_{F}$ in $K_{\hat{F}}^{\ast}$ and let $x_{\hat{F}}$ be the projection of $x$ to $K_{\hat{F}}^{\ast}$. Then $V_{\hat{F}}$ is homeomorphic to the complement of a point of degree $2$ in a finite connected sum of $2$-connected graphs. To see this, view $V_{F}$ as the complement of a point in the circle $S^{1}$ and then perform the connected sum operations required to obtain $K_{\hat{F}}^{\ast}$ from $K_{F}^{\ast}$. It follows then from Lemma 4.6 and Remark 4.7 that $V_{\hat{F}}\setminus\{x_{\hat{F}}\}$ has at most two components. But $V\setminus\{x\}$ is the ascending union of the preimages of the $V_{\hat{F}}\setminus\{x_{\hat{F}}\}\subset K_{\hat{F}}^{\ast}$ as $\hat{F}$ ranges over all finite b-subtrees containing $F$. Hence $V$ itself has at most two components. This establishes the forward implication.
@@
To prove the reverse implication, assume that no such $F$ exists as in the statement. Then the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ all have $2$ components for a cofinal chain of finite b-subtrees $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ as above. As argued above, the inclusion of the preimage of $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ in $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ induces a bijection on components and so. Thus the preimages of the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ in $Z$ express $Z\setminus\{x,y\}$ as a disjoint union of two ascending unions of connected open subspaces. It follows that $Z\setminus\{x,y\}$ has two components.
The last sentence of the statement follows from the same argument given above showing that $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ has no more components than $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$. ∎
@@
We will define $\tilde{v}$ as the limit of a system of lifts $v_{F}\in K_{F}^{\ast}$ of $v$ where $F\supset F_{0}$. These lifts need to be consistent with respect to the bonding maps so it suffices to show how to lift $v_{F}$ to $v_{\hat{F}}$ for finite b-subtrees $\hat{F}\supset F\supset F_{0}$ for which no b-subtree lies properly between $\hat{F}$ and $F$. Then the reduced partial union $K_{\hat{F}}^{\ast}$ is obtained from $K_{F}^{\ast}$ by performing a connected sum operation with a $\theta$-graph $K_{t}^{\ast}$. In the case that the connected sum operation is performed away from $v_{F}$ then we let $v_{\hat{F}}=v_{F}$ in $K_{\hat{F}}^{\ast}$. In the case that the connected sum operation is performed at $v_{F}$, we let $v_{\hat{F}}$ be the essential vertex of $K_{t}^{\ast}$ that is not involved in the connected sum operation. Notice that, in either case, $v_{\hat{F}}$ is the unique vertex of $K_{\hat{F}}^{\ast}$ of degree $\deg(v_{F})$ that projects onto $v_{F}$ in $K_{F}^{\ast}$. Thus the limit $\tilde{v}\in Z$ of the $v_{F}$ is the unique point in $Z$ projecting to $v$ in $K_{F_{0}}^{\ast}$ and projecting to an essential vertex of degree $\deg(v_{F_{0}})$ in every $K_{F}^{\ast}$ with $F\supset F$.

### v2

@@ 5.1. Proof of the impliciation (1$\Rightarrow$(3) of Theorem A
Let $G$ be a hyperbolic group and assume that $G$ is not virtually free (in particular, it is not finite and not virtually cyclic). Then $G$ has a $1$-ended quasi-convex subgroup.
@@
By the Dunwoody’s Accessibility Theorem [Dun85], there is a splitting of $G$ as a finite graph of groups whose edge groups are finite and whose vertex groups have at most $1$ end. Since graph of finite groups always yields a virtually free group (see Proposition 11 in [Ser03]), at least one of the vertex groups above must be 1-ended. Finally, this vertex group is quasi-convex e.g. by the discussion in the paragraph right before Corollary 3.7. ∎
Next rather well known result follows e.g. by [KB02, Theorem 2.28].
@@
We are now ready to prove the main result of this section: the impliciation (1$\Rightarrow$(3) of Theorem A. We restate and prove the implication below.
@@
For the sake of reaching a contradiction, assume that $\partial G$ is a tree of graphs and $G$ has a rigid cluster factor $H$ that is not virtually free. Then, by Lemma 5.9, the rigid cluster factor $H$ has a $1$-ended quasi-convex subgroup $K$. Then, since $H$ is quasi-convex in $G$, the boundary $\partial K$ is a closed subspace of $\partial G$. Since $K$ is $1$-ended, the boundary $\partial K$ is connected. By Lemma 5.10, the boundary $\partial K$ has at least two points. Moreover, the boundary $\partial G$ is connected, locally connected and has no cutpoints [KB02, Proposition 7.1, Theorem 7.2]. Thus we can apply Proposition 5.7 to the subspace $\partial K\subset\partial G$ to conclude that $\partial K$ has uncountable intersection with some subnecklace $N^{\prime}\subset\partial G$. Recalling Remark 5.2, we see that $\partial K$ has uncountable intersection with a necklace class $N$ of $\partial G$. Then, since $\partial K\subset\partial H$, we see that $\partial H$ has uncountable intersection with $N$. This contradicts Lemma 5.8. ∎
@@ 5.2. Twin graphs and trees of thick $\theta$-graphs
Recall that a $\theta$-graph is a finite graph with two vertices, at least one edge and no loop edges. Note that a $\theta$-graph is 2-connected if it has at least two edges. We say that a $\theta$-graph is thick if it has at least three edges. In this subsection we will study connectedness properties of finite connected sums of $\theta$-graphs which we call twin graphs and characerize in terms of their Bowditch cut pairs. We will use these results in the following subsection to study trees of thick $\theta$-graphs and to prove the implication (1$\Rightarrow$(3) of Theorem B.
@@
Lemma 5.15 (Bowditch [Bow98, Lemma 3.8]).
@@
We now prove (2). Suppose first that $\deg(y_{i})\neq\bigl{|}\pi_{0}\bigl{(}X_{i}\setminus\{x_{i},y_{i}\}\bigr{)}\bigr{|}$. Without loss of generality, we can assume $i=1$. Notice that $\deg(y_{1})$ is equal to the cardinality of the blow-up divisor $P_{y_{1}}$ and that $\bigl{|}\pi_{0}\bigl{(}X_{1}\setminus\{x_{1},y_{1}\}\bigr{)}\bigr{|}$ is equal to the number of components of the blow-up $X_{1}^{\#}(\{x_{1},y_{1}\})$. Then, by Lemma 4.6, we have distinct $p,p^{\prime}\in P_{y_{1}}$ that are contained in the same component $C$ of $X_{1}^{\#}(\{x_{1},y_{1}\})$. But $X_{1}^{\#}(x_{1})$ is a subspace of $X$ with $y_{1}\in X_{1}^{\#}(x_{1})$ and $y_{2}\notin X_{1}^{\#}(x_{1})$ so $X_{1}^{\#}(\{x_{1},y_{1}\})$ is a subspace of $X^{\#}(\{y_{1},y_{2}\})$. Then $C$ is a connected subspace of $X^{\#}(\{y_{1},y_{2}\})$ containing $p,p^{\prime}\in P_{y_{1}}$, which by Remark 5.13, implies that $y_{1}$ and $y_{2}$ are not twins.
Suppose now that $\ell(\sigma_{1})\neq\sigma_{2}$. Without loss of generality (by swapping, if necessary, the indices $i\in\{1,2\}$) there are $p,p^{\prime}\in P_{x_{1}}$ contained in the same component $C_{1}$ of $X_{1}^{\#}(\{x_{1},y_{1}\})$ such that $\ell(p)$ and $\ell(p^{\prime})$ are contained in distinct components $C_{2}$ and $C_{2}^{\prime}$ of $X_{2}^{\#}(\{x_{2},y_{2}\})$. By Lemma 4.6, there are $p_{2},p_{2}^{\prime}\in P_{y_{2}}$ with $p_{2}\in C_{2}$ and $p_{2}^{\prime}\in C_{2}^{\prime}$. We can view $X^{\#}(\{y_{1},y_{2}\})$ as the connected sum of $X_{1}^{\#}(y_{1})$ and $X_{2}^{\#}(y_{2})$ at $x_{1}$ and $x_{2}$, along $\ell\colon P_{x_{1}}\to P_{x_{2}}$. Then $C_{1}\cup C_{2}\cup C_{2}^{\prime}$ is a connected subspace of $X^{\#}(\{y_{1},y_{2}\})$ containing both $p_{2}$ and $p_{2}^{\prime}$ which, by Remark 5.13, implies that $y_{1}$ and $y_{2}$ are not twins. Thus we have proved the forward implication of (2).
@@
Lemma 5.22 (Bowditch [Bow98, Lemma 3.18]).
@@
For an essential twin pair $\{x,\bar{x}\}$, let $n_{\{x,\bar{x}\}}$ be the maximal number of essential vertices contained in a component of $X\setminus\{x,\bar{x}\}$. The complexity of $\{x,\bar{x}\}$ is the nonnegative integer $c_{\{x,\bar{x}\}}=n-n_{\{x,\bar{x}\}}-2$, where $n$ is the number of essential vertices of $X$. Note that $c_{\{x,\bar{x}\}}$ is the number of essential vertices in all but a “largest” component of $X\setminus\{x,\bar{x}\}$. So it suffices to find an essential twin pair $\{x,\bar{x}\}$ for which $c_{\{x,\bar{x}\}}=0$. This is the base case of an induction starting with an arbitrary essential twin pair $\{x,\bar{x}\}$. It suffices to show that if $c_{\{x,\bar{x}\}}>0$ then we can find an essential twin pair $\{y,\bar{y}\}$ for which $c_{\{y,\bar{y}\}}<c_{\{x,\bar{x}\}}$. Let $C^{\prime}$ be a component of $X\setminus\{x,\bar{x}\}$ containing $n_{\{x,\bar{x}\}}$ essential vertices. Since $n_{\{x,\bar{x}\}}=n-c_{\{x,\bar{x}\}}-2<n-2$, some other component $C$ of $X\setminus\{x,\bar{x}\}$ contains an essential vertex $y$. We will show that $c_{\{y,\bar{y}\}}<c_{\{x,\bar{x}\}}$.
@@
Suppose $C$ contains at least one essential vertex. By Lemma 4.6 there is a topological edge $e_{x}$ with endpoint $x$ and whose interior is contained in $C$. Similarly, there is a topological edge $e_{\bar{x}}$ with endpoint $\bar{x}$ and whose interior is contained in $C$. Let $p_{x}$ be a point in the interior of $e_{x}$ and let $p_{\bar{x}}$ be a point in the interior of $e_{\bar{x}}$. Then $X^{\#}(\{p_{x},p_{\bar{x}}\})$ has two components, one of which contains $x$ and $\bar{x}$ and the other of which is naturally a subspace of $C$. Gluing together the pairs of points of the blow-up divisors $P_{p_{x}}\cup P_{p_{\bar{x}}}\subset X^{\#}(\{p_{x},p_{\bar{x}}\})$ that are contained in the same component results in a disjoint union of two graphs: a thick $\theta$-graph $X_{1}$ with essential vertices $x$ and $\bar{x}$ and a graph $X_{2}$ with two less essential vertices than $X$. Moreover, the original graph $X$ is a connected sum of $X_{1}$ and $X_{2}$. Thus, by induction, it suffices to show that $X_{2}$ is a twin graph. This follows immediately from Proposition 5.19(1). ∎
@@
In the inductive step, we can write $X$ as a connected sum $X_{1}\#X_{2}$ where $X_{1}$ is a twin graph and $X_{2}$ is a thick $\theta$-graph and the connected sum is performed at points $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Since performing a connected sum with a $\theta$-graph at an essential vertex does not change the topology, we can assume that $\deg(x_{1})=\deg(x_{2})=2$. Up to homeomorphism, we can assume that $x$ and $y$ are not contained in the blow-up divisor $P_{x_{1}}=P_{x_{2}}\subset X_{1}\#X_{2}$. In fact, we can assume that $x$ and $y$ are not contained in $e\setminus\{v,\bar{v},x_{2}\}\subset X_{1}\#X_{2}$ where $e$ is the topological edge of $X_{2}$ containing $x_{2}$ and $\{v,\bar{v}\}$ is the unique essential twin pair of $X_{2}$. We consider first the case where one of the points, say $x$, is contained in $X_{2}\setminus e\subset X_{1}\#X_{2}$. Then $y$ cannot be contained in the topological edge of $X_{2}$ that contains $x$ since otherwise $X\setminus\{x,y\}$ would be disconnected. Hence $x$ and $y$ are contained in distinct components of the complement of $\{v,\bar{v}\}$, which is a twin pair of $X$ by Proposition 5.19(1). It remains then to consider the case where $x,y\in X_{1}\setminus\{x_{1}\}\subset X_{1}\#X_{2}$. By Proposition 5.19(1), the degree $2$ points $x$ and $y$ are not twins in $X_{1}$ and so $X_{1}\setminus\{x,y\}$ is connected. Then, by induction, there exists an essential twin pair $\{w,\bar{w}\}$ of $X_{1}$ separating $x$ from $y$. By Proposition 5.19(1), the pair $\{w,\bar{w}\}$ continues to be a twin pair in $X_{1}\#X_{2}$. Since the connected sum operation with the $\theta$-graph $X_{2}$ preserves components of the complement of $\{w,\bar{w}\}$ we see that $\{w,\bar{w}\}$ continues to separate $x$ and $y$ in $X=X_{1}\#X_{2}$. ∎
5.3. Proof of the impliciation (1$\Rightarrow$(3) of Theorem B
We now turn our attention from twin graphs towards trees of thick $\theta$-graphs in order to prove the implication (1$\Rightarrow$(3) of Theorem B, i.e., we will show that if $G$ is a $1$-ended hyperbolic group that is not cocompact Fuchsian and $\partial G$ is a tree of thick $\theta$-graphs then $G$ has no rigid JSJ factors.
Recall from Section 3.1 that the degree of a point $p$ in a compact space $X$ is the number of ends of the locally compact space $X\setminus\{p\}$. Since $X$ is compact, any compact subset of $X\setminus\{p\}$ is the complement of some open neighborhood $U$ of $p$. If $U$ is also connected then the degree of $p$ is at least the number of components of $U\setminus\{p\}$ since, in this case, each component of $U\setminus\{p\}$ contains at least one end of $X\setminus\{p\}$. It follows that, if $p$ has a basis $\{U_{\alpha}\}_{\alpha}$ of connected open neighborhoods then the degree of $p$ is the supremum of the number of components of the $U_{\alpha}\setminus\{p\}$.
@@ 5.3. Proof of the impliciation (1$\Rightarrow$(3) of Theorem B
We first prove the contrapositive of the forward implication. Assume that for any finite b-subtree $F_{0}\subset T$ there exists a finite b-subtree $F\supset F_{0}$ such that the projection $x_{F}$ of $x$ to the reduced partial union $K_{F}^{\ast}$ has degree less than $3$. Let $U$ be an open neighborhood of $x$ in $Z$. Then there exists a reduced partial union $K_{F}^{\ast}$ and an open neighborhood $V_{F}$ of the projection $x_{F}$ of $x$ in $K_{F}^{\ast}$ such that the preimage of $V_{F}$ in $Z$ is contained in $U$. By our assumption, we can choose $K_{F}^{\ast}$ so that $x_{F}$ has degree $2$ and so we can choose $V_{F}$ so that it is homeomorphic to an open interval. By Lemma 4.13, the preimage $V$ of $V_{F}$ in $Z$ is connected. By the paragraph preceding the statement of the lemma, it suffices to show that $V\setminus\{x\}$ has at most two components in order to establish that $x$ has degree at most $2$. Now, let $K_{\hat{F}}^{\ast}$ be any reduced partial union with $\hat{F}\supset F$, let $V_{\hat{F}}$ be the preimage of $V_{F}$ in $K_{\hat{F}}^{\ast}$ and let $x_{\hat{F}}$ be the projection of $x$ to $K_{\hat{F}}^{\ast}$. Then $V_{\hat{F}}$ is homeomorphic to the complement of a point of degree $2$ in a finite connected sum of $2$-connected graphs. To see this, view $V_{F}$ as the complement of a point in the circle $S^{1}$ and then perform the connected sum operations required to obtain $K_{\hat{F}}^{\ast}$ from $K_{F}^{\ast}$. It follows then from Lemma 4.6 and Remark 4.7 that $V_{\hat{F}}\setminus\{x_{\hat{F}}\}$ has at most two components. But $V\setminus\{x\}$ is the ascending union of the preimages of the $V_{\hat{F}}\setminus\{x_{\hat{F}}\}\subset K_{\hat{F}}^{\ast}$ as $\hat{F}$ ranges over all finite b-subtrees containing $F$. Hence $V$ itself has at most two components. This establishes the forward implication.
@@
To prove the reverse implication, assume that no such $F$ exists as in the statement. Then the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ all have $2$ components for a cofinal chain of finite b-subtrees $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ as above. As argued above, the inclusion of the preimage of $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ in $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ induces a bijection on components. Thus the preimages of the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ in $Z$ express $Z\setminus\{x,y\}$ as a disjoint union of two ascending unions of connected open subspaces. It follows that $Z\setminus\{x,y\}$ has two components.
The last sentence of the statement follows by the same argument given above showing that $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ has no more components than $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$. ∎
@@
We will define $\tilde{v}$ as the limit of a system of lifts $v_{F}\in K_{F}^{\ast}$ of $v$ where $F\supset F_{0}$. These lifts need to be consistent with respect to the bonding maps so it suffices to show how to lift $v_{F}$ to $v_{\hat{F}}$ for finite b-subtrees $\hat{F}\supset F\supset F_{0}$ for which no b-subtree lies properly between $\hat{F}$ and $F$. Then the reduced partial union $K_{\hat{F}}^{\ast}$ is obtained from $K_{F}^{\ast}$ by performing a connected sum operation with a $\theta$-graph $K_{t}^{\ast}$. In the case that the connected sum operation is performed away from $v_{F}$ then we let $v_{\hat{F}}=v_{F}$ in $K_{\hat{F}}^{\ast}$. In the case that the connected sum operation is performed at $v_{F}$, we let $v_{\hat{F}}$ be the essential vertex of $K_{t}^{\ast}$ that is not involved in the connected sum operation. Notice that, in either case, $v_{\hat{F}}$ is the unique vertex of $K_{\hat{F}}^{\ast}$ of degree $\deg(v_{F})$ that projects onto $v_{F}$ in $K_{F}^{\ast}$. Thus the limit $\tilde{v}\in Z$ of the $v_{F}$ is the unique point in $Z$ projecting to $v$ in $K_{F_{0}}^{\ast}$ and projecting to an essential vertex of degree $\deg(v_{F_{0}})$ in every $K_{F}^{\ast}$ with $F\supset F_{0}$.